# Polynomial Time Convergence of the Iterative Evaluation of Datalogo Programs

### 21 Dec 2023 v1

@@ 3 Small Support Representations
To streamline our analysis, we will assume that $\lambda\geq 2$. Otherwise, we can pretend that this is the case by adding an unused terminal to $\Sigma$. Then, we will only need to increment the value of $\lambda$ by one in the final bound.
Consider an arbitrary word $w$ in our language $L$. Let $\mathcal{L}$ be a linear set that contains $w$. Since $\mathcal{L}$ is linear it can be defined as follows:
$\mathcal{L}=$$\{{\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{\ell}\bm{v}_{\ell}}\mid{k_{1},\ldots,k_{\ell}\in\mathbb{N}}\}$
For any word $w$, by renumbering the basis vectors we can assume that basis vectors $\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{h}$ have non-zero coefficients $k_{1},k_{2},\ldots,k_{h}>0$. So we have:
$w=$ $\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{h}\bm{v}_{h}$
@@
$\sum_{i\in H_{1}}\bm{v}_{i}=\sum_{i\in H_{2}}\bm{v}_{i}$ (17)
We will prove this claim by upper-bounding
$K:=|\{\sum_{i\in H}\bm{v}_{i}\;|\;H\subseteq[h]\}|,$
the number of distinct vectors we can generate by choosing a subset of vectors from $\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{h}$ (we choose each basis vector up to one copy).
Let $M$ be the maximum 1-norm of any $\bm{v}_{i}$, i.e., $||\bm{v}_{i}||_{1}\leq M$, for $i\in[h]$. Thus, we have $||\sum_{i\in H}\bm{v}_{i}||_{1}\leq Mh$ for any $H\subseteq[h]$. We use the following well-known fact: the number of distinct vectors in $Z^{d}_{\geq 0}$ with 1-norm of $k$ is exactly (${k+d-1\choose d-1}\leq(k+d-1)^{d-1}$. In our case, $k\leq Mh$ and $d=\sigma$. Thus, we have $K\leq(Mh+\sigma-1)^{\sigma-1}(Mh+1)\leq(Mh+\sigma)^{\sigma}$. If $\sigma=1$ (there exits only one terminal), we have a tighter bound of $K\leq M+(M-1)+\ldots+M-(h-1)=h(2M-(h-1))/2$.
We aim to show
$2^{h}>M,$ (18)
which will imply the existence of a desired pair of $H_{1}$ and $H_{2}$ satisfying Eqn. (17) due to the pigeon hole principle.
We first consider the case when $\sigma\geq 2$. We shall now establish
$2^{h}>(Mh+\sigma)^{\sigma}\textnormal{ when $\sigma\geq 2$}$ (19)
By taking the logarithm, the inequality (19) is equivalent to:
$h>\sigma\lg(Mh+\sigma)$ (20)
We know from our main Theorem 1.2 that $M\leq\lambda^{n(n+3)/2}$. Thus we know that the following equation would imply Eqn. (20):
$h>\sigma\lg(h\lambda^{n(n+3)/2}+\sigma)$ (21)
Since $\lg x$ is sub-additive when $x\geq 2$, using the assumptions that $\lambda,\sigma\geq 2$, it is sufficient to show:
$\;\;h>\sigma\lg(h\lambda^{n(n+3)/2})+\sigma\lg\sigma=\sigma\lg h+\sigma(n(n+3)/2)\lg\lambda+\sigma\lg\sigma$
$\Leftrightarrow$ $\;\;h-\sigma\lg h>\sigma(n(n+3)/2)\lg\lambda+\sigma\lg\sigma$
We show that
$h-\sigma\lg h\geq h/2$
when $h\geq 8\sigma\lg\sigma$.
Since $h/2-\sigma\lg h$ is increasing in $h$ when $h\geq 8\sigma\lg\sigma$, we have $h/2-\sigma\lg h\geq 4\sigma\lg\sigma-\sigma\lg(8\sigma\lg\sigma)=\sigma\lg(\sigma^{3}/(8\lg\sigma))\geq 0$ when $\sigma\geq 2$, as desired. Thus, we have shown that $2^{h}>K$ when $\sigma\geq 2$.
If $\sigma=1$, as mentioned before, we have $K\leq M+(M-1)+\ldots+M-(h-1)=h(2M-(h-1))/2$. Thus, we have:
$2^{h}>K$
[…]
which is true due to the assumption that $h>2(\sigma(n(n+3)/2)\lg\lambda+4\sigma\lg\sigma)=n(n+3)\lg\lambda$; recall $\sigma=1$. Thus, for all values of $\sigma$, we have shown that $2^{h}>K$, which implies the existence of $H_{1}\neq H_{2}\subseteq[h]$ satisfying Eqn. (17).
Now that we have established the existence of a desired pair of $H_{1}$ and $H_{2}$, we now explain how to construct a new representation of $w$ that contains less basis vectors. First observe that one of two sets $H_{1},H_{2}$ doesn’t contain the other since no basis vectors are $\bm{0}$ and we have Eqn. (17). Let $i$ be $\arg\min_{i^{\prime}\in H_{1}-H_{2}}k_{i^{\prime}}$, breaking ties arbitrarily. Then for $j\in H_{1}-H_{2}$ let $k_{j}^{\prime}=k_{j}-k_{i}$, for $j\in H_{2}-H_{1}$ let $k_{j}^{\prime}=k_{j}+k_{i}$, and for all other $i$ let $k_{j}^{\prime}=k_{j}$.
@@
$w=$ $\;\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{h}\bm{v}_{h}$
$=$ $\;\bm{v}_{0}+k^{\prime}_{1}\bm{v}_{1}+\ldots+k^{\prime}_{h}\bm{v}_{h}$
However, this new representation $\langle k^{\prime}_{1},k^{\prime}_{2},\ldots,k^{\prime}_{h}\rangle$ has a strictly smaller support since $k^{\prime}_{i}=0$. This contradicts the minimality in the definition of $h$.
Thus, we have shown that $w$ has a linear representation of support size at most $2(\sigma(n(n+3)/2)\lg\lambda+4\sigma\lg\sigma)$ assuming $\lambda\geq 2$. To remove this assumption we add an unused terminal to $\Sigma$, which increments the value of $\lambda$ by one in the bound. ∎
4 Bounding the Number of Iterations
This section is devoted to proving Theorem 1.1.
Consider an arbitrary IDB variable $X_{i}$ and let $L$ be the CFL associated with this variable. For the sake of analysis we impose a linear order on the words $w\in L$. Let $\mathcal{M}$ be a semi-linear set that satisfies the properties stated in Theorem 1.2. Let $\ell(w)$ be defined as the minimum 1-norm of all possible linear representations of $w$ in terms of the offset and basis vectors of some $\mathcal{L}\in\mathcal{M}$. We order the words $w$ by increasing order of $\ell(w)$ breaking ties arbitrarily. We will show adding a summand $\Psi(w)$ with $\ell(w)>ph$ doesn’t change the following summation,
$\sum_{\ell=1}^{\ell(w)-1}\sum_{w^{\prime}:\ell(w^{\prime})=\ell}\Psi(w^{\prime}),$ (22)
@@ 4 Bounding the Number of Iterations
$h=2(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma)$
Consider an arbitrary $w\in L$ with $\ell(w)>ph$. We know from Lemma 1.3 that there is a sparse linear representation
$\Psi(w)=\bm{v}_{0}+k_{1}\bm{v}_{\gamma(1)}+\ldots+k_{h}\bm{v}_{\gamma(h)}$
of the Parikh image of $w$ in terms of the offset and basis vectors for some linear set $\mathcal{L}\in\mathcal{M}$. Since $\ell(w)>ph$, by the pigeon hole principle there must exist an $j\in[h]$ such that $k_{j}\geq p+1$. We can assume without loss of generality (by renumbering the basis vectors) that $j=1$. Now consider the Parikh images:
$\bm{v}_{0}+(k_{1}-g)\bm{v}_{\gamma(1)}+\ldots+k_{h}\bm{v}_{\gamma(h)}$
for $g\in[1,p+1]$. By the first property of Theorem 1.2 there are $p$ words $w_{1},\ldots,w_{p}$ in $L$ that have these Parikh images. Further, since for all $i\in[1,p+1]$ $\ell(w_{i})<\ell(w)$ values, each $\Psi(w_{i})$ appears in Eqn. (22). But then by the $p$-stability of the element corresponding to $\bm{v}_{\gamma(1)}$, adding the summand $\Psi(w)$ doesn’t change the sum.
Then by condition 2(d) of Theorem 1.2, we know that any word $w\in L$ with $\ell(w)\leq ph$ can be yielded by a parse tree of depth at most
$(ph+1)n(n+3)/2$
which substituting in the value of $h$ is upper bounded by
$q:=pn(n+3)\cdot(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma+1)$ (23)
Thus, we have shown that :
$f_{i}^{(q)}(\bm{0})=f_{i}^{(q+1)}(\bm{0})$
where
$f_{i}^{(q)}(\bm{0})=\sum_{T\in\mathcal{T}^{i}_{q}}Z(T).$
and $\mathcal{T}^{i}_{q}$ denotes the collection of the parse trees of depth at most $q$ starting with $X_{i}$. This completes the proof.
5 Related Work
If the semiring is naturally ordered111 $\bm{S}$ is naturally ordered if the relation $x\preceq_{S}y$ defined as $\exists z:x\oplus z=y$ is a partial order. , then the least fixed point of a $\text{\sf Datalog}^{\circ}$ program is the least fixed point of $f$ under the same partial order extended to $S^{n}$ componentwise. This is the least fixed point semantics of a $\text{\sf Datalog}^{\circ}$ program. Under (ground-atom) set containment, this function has a unique least fixed point which can be obtained via repeatedly applying the natural iterative evaluation algorithm that applies ICO until a fixed point is reached [1]. This least fixed point is the semantics of the given Datalog program.
Computing the least fixed point solution to a recursive $\text{\sf Datalog}^{\circ}$ program boils down to solving fixed point equations over semirings. In particular, we are given a multi-valued polynomial function $f:S^{n}\to S^{n}$ over a commutative semiring, and the problem is to compute a (pre-) fixpoint of $f$, i.e. a point $x\in S^{n}$ where $x=f(x)$. As surveyed in much details in [10], this problem was studied in a very wide range of communities, such as in automata theory [14], program analysis [4, 18], graph algorithms [3, 17, 16] communities since the 1970s. (See [21, 8, 15, 7, 23] and references thereof).
The first approach is to keep iterating until a fixpoint is reached; in different contexts, this has different names: the naive algorithm, Jacobi iteration, Gauss-Seidel iteration, or Kleene iteration. The main advantage of this approach, which is studied in this paper, is that it assumes less about the underlying algebraic structure: we do not need both left and right distributive law, and do not need to assume a closure operator. The second approach is based on Gaussian elimination (Gauss-Jordan elimination), which, assuming oracle access to the solution $x^{*}$ of the 1D problem $x=1+ax$, can solve the algebraic path problem in $O(n^{3})$-time [21, 15]. The third approach is based on specifying the solutions based on the free semiring generated when viewing $\bm{A}$ as the adjacency matrix of a graph [22]. The underlying graph structure (such as planarity) may be exploited for very efficient algorithms [17, 16].
We also discuss the previous work on the fixed point solution to $x=f(x)$ in the non-linear case. See [20, 19, 13] for the study when $f$’s component functions are multivariate polynomials over the Kleene algebra. In the non-linear case, Kleene iteration may not converge in a finite number of steps. Thus, methods based on Galois connection or on widening/narrowing approaches [5] were proposed. These approaches are (discrete) lattice-theoretic. More recently, a completely different approach drawing inspiration from Newton’s method for solving a system of (real) equations was proposed [8, 6]. Newton’s method converges at least as fast as Kleene iteration, but each iteration could be more costly it need to compute a matrix inverse [6]. Recently, semi-naïve evaluation has been extended for a higher-order functional language called Datafun [2]. The book [7] contains many fundamental results on algebraic structures related to semirings and computational problems on such structures.
There are a daunting number of different statements, proofs and extensions of Parikh’s theorem in the literature. See for example [11]. As far as we know our extension of Parikh’s theorem does not appear in the literature, but given the volume of the work in this area, it is hard to be totally confident about the novelty of our extension.
Appendix A Semirings
In this section, we introduce the notation and terminology used throughout the paper.
Definition A.1 (Semiring).
A semiring [7] is a tuple $\bm{S}=(S,\oplus,\otimes,0,1)$ where
  • $\oplus$ and $\otimes$ are binary operators on $S$,
  • $(S,\oplus,0)$ is a commutative monoid, meaning $\oplus$ is commutative and associative, and 00 is the identity for $\oplus$,
  • $(S,\otimes,1)$ is a monoid, meaning $\otimes$ is associative, and 00 is the identity for $\oplus$,
  • 00 annilates every element $a\in S$, that is $a\otimes 0=0\otimes a=0$, and
  • $\otimes$ distributes over $\oplus$.
A commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$ is a semiring where additionally $\otimes$ is commutative.
In the following fact about $p$-stable commutative semiring will be useful.
Proposition A.2.
Given a $p$-stable commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$, for any $u\in S$, we have $pu=(p+1)u$, where $pu$ here is shorthand for $\oplus_{i=1}^{p}u$.
Proof.
This follows directly from the $p$-stability of 1, and the fact that $11^{2}=$ 1. ∎

### 21 Feb 2024 v2

@@ 5. Small Support Representations
To streamline our analysis, we will assume that $\lambda\geq 2$. This is without loss of generality because in the case that $\lambda=1$ we can add an unused terminal to $\Sigma$. The value of $\lambda$ will increase by one in the final bound for this boundary case. For an arbitrary word $w$ in our language $L$, suppose we are given a linear representation of $\Psi(w)$,
$\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{m}\bm{v}_{m},$
where
$k_{1},k_{2},\ldots k_{m}>0\textnormal{ and }$
$m>h:=2(\sigma(n(n+3)/2)\lg\lambda+4\sigma\lg\sigma).$
It suffices to find another linear representation of $\Psi(w)$,
$\bm{v}_{0}+k^{\prime}_{1}\bm{v}_{1}+\ldots+k^{\prime}_{m}\bm{v}_{m}$
such that
(18) $k_{1}+k_{2}+\ldots+k_{m}=k^{\prime}_{1}+k^{\prime}_{2}+\ldots+k^{\prime}_{m}\textnormal{ and }$
(19) $k^{\prime}_{i}=0\textnormal{ for some }i\in[m]:=\{1,2,\ldots,m\}.$
A key step to our proof is showing that there exist two distinct subsets $H_{1}$ and $H_{2}$ of $[m]:=\{1,2,\ldots,m\}$ such that
(20) $\sum_{i\in H_{1}}\bm{v}_{i}=\sum_{i\in H_{2}}\bm{v}_{i}$
We will prove this claim by proving that
(21) $K$ :=|$:=|\{\sum_{i\in H}\bm{v}_{i}\;|\;\emptyset\neq H\subseteq[m]\}|<2^{m}.$
Note that $K$ is the number of distinct vectors we can generate by summing a non-empty subset of vectors from $\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{m}$. The existence of the desired pair of $H_{1}$ and $H_{2}$ satisfying (20) will then follow from pigeonhole principle.
Let $M$ be the maximum 1-norm of any $\bm{v}_{i}$, i.e., $||\bm{v}_{i}||_{1}\leq M$, for all $i\in[m]$. Thus, we have $||\sum_{i\in H}\bm{v}_{i}||_{1}\leq Mm$ for any $H\subseteq[m]$. We use the following well-known fact: the number of distinct vectors in $\mathbb{N}^{d}$ with 1-norm of $k$ is exactly $\binom{k+d-1}{d-1}\leq(k+d-1)^{d-1}$ (see [14]). In our case, $k\leq Mm$ and $d=\sigma$. Thus, we have $K\leq(Mm+\sigma-1)^{\sigma-1}(Mm+1)\leq(Mm+\sigma)^{\sigma}$. If $\sigma=1$ (there exists only one terminal), we have a tighter bound of $K\leq M+(M-1)+\ldots+M-(m-1)=m(2M-(m-1))/2$.
To prove (21), it remains to show that $(Mm+\sigma)^{\sigma}<2^{m}$ when $\sigma\geq 2$ and that $h(2M-m+1)/2<2^{m}$ when $\sigma=1$. We consider two cases: $\sigma\geq 2$ and $\sigma=1$.
Case i: $\sigma\geq 2$. We shall now establish
(22) $2^{m}>(Mm+\sigma)^{\sigma}\textnormal{ when $\sigma\geq 2$}$
By taking the logarithm, the inequality (22) is equivalent to:
(23) $m>\sigma\lg(Mm+\sigma)$
We know from Theorem 3.1 (c) that $M\leq\lambda^{n(n+3)/2}$. Thus we know that the following equation would imply Eqn. (23):
(24) $m>\sigma\lg(m\lambda^{n(n+3)/2}+\sigma)$
Using the fact that $\lg x$ is sub-additive when $x\geq 2$ and the assumptions that $\lambda,\sigma\geq 2$, we have
(25) $\sigma\lg(m\lambda^{n(n+3)/2})+\sigma\lg\sigma>\sigma\lg(m\lambda^{n(n+3)/2}+\sigma)$
Thus, it is sufficient to show:
$\;\;m>\sigma\lg(m\lambda^{n(n+3)/2})+\sigma\lg\sigma$
$\;\;=\sigma\lg m+\sigma(n(n+3)/2)\lg\lambda+\sigma\lg\sigma$
(26) $\Leftrightarrow$ $\;\;m-\sigma\lg m>\sigma(n(n+3)/2)\lg\lambda+\sigma\lg\sigma$
We show that
$m-\sigma\lg m\geq m/2$
when $m\geq 8\sigma\lg\sigma$: Since $m/2-\sigma\lg m$ is increasing in $m$ when $m\geq 8\sigma\lg\sigma$, we have $m/2-\sigma\lg m\geq 4\sigma\lg\sigma-\sigma\lg(8\sigma\lg\sigma)=\sigma\lg(\sigma^{3}/(8\lg\sigma))\geq 0$ when $\sigma\geq 2$, as desired.
Thus, we have
(27) $m-\sigma\lg m\geq m/2>\sigma(n(n+3)/2)\lg\lambda+4\sigma\lg\sigma,$
where the second inequality follows from the fact that $m>h$. From Eqn. (22), (23), (24), (25), (26), and (27) we have $2^{m}>K$ when $\sigma\geq 2$.
Case ii: $\sigma=1$. If $\sigma=1$, as mentioned above, we have
$K$ $\leq M+(M-1)+\ldots+M-(m-1)$
$=m(2M-(h-1))/2$
$<Mm$
@@
$\leq\lambda^{n(n+3)/2}2^{m/2}$
$\leq 2^{m}.$
The last inequality is true due to the assumption that $\sigma=1$ and
$m$ $>h=2(\sigma(n(n+3)/2)\lg\lambda+4\sigma\lg\sigma)\geq n(n+3)\lg\lambda$
Thus, we have shown that $2^{m}>K$ for all $\sigma\geq 1$, which establishes the existence of $H_{1}\neq H_{2}\subseteq[m]$ satisfying Eqn. (20).
We now explain how to construct a new representation of $\Psi(w)$ that contains less basis vectors. First observe that one of two sets $H_{1},H_{2}$ doesn’t contain the other since no basis vectors are $\bm{0}$ and we have Eqn. (20). Let $i$ be $\arg\min_{i^{\prime}\in H_{1}-H_{2}}k_{i^{\prime}}$, breaking ties arbitrarily. Then for $j\in H_{1}-H_{2}$ let $k_{j}^{\prime}=k_{j}-k_{i}$, for $j\in H_{2}-H_{1}$ let $k_{j}^{\prime}=k_{j}+k_{i}$, and for all other $i$ let $k_{j}^{\prime}=k_{j}$.
@@
$w=$ $\;\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{m}\bm{v}_{m}$
(28) $=$ $\;\bm{v}_{0}+k^{\prime}_{1}\bm{v}_{1}+\ldots+k^{\prime}_{m}\bm{v}_{m}$
However, this new representation $\langle k^{\prime}_{1},k^{\prime}_{2},\ldots,k^{\prime}_{m}\rangle$ has a strictly smaller support since $k^{\prime}_{i}=0$. Observe that $k_{1}+k_{2}+\ldots+k_{m}=k^{\prime}_{1}+k^{\prime}_{2}+\ldots+k^{\prime}_{m}$. Thus, we have found another linear representation $\bm{v}_{0}+k^{\prime}_{1}\bm{v}_{1}+\ldots+k^{\prime}_{m}\bm{v}_{m}$ of $\Psi(w)$ that has a smaller support than the given linear representation $\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{m}\bm{v}_{m}$ preserving the 1-norm value in the linear representation.
We can repeat this process until we obtain a linear representation of support size at most $h$. Finally, recall that we assumed $\lambda\geq 2$. To remove this assumption, as mentioned at the beginning, we can add an unused terminal to $\Sigma$, which increments the value of $\lambda$ by one in the bound. ∎
6. Bounding the Number of Iterations
This section is devoted to proving Theorem 1.1, restated here.
Theorem 6.1 (Theorem 1.1 Restated ).
Let $\bm{S}$ be a $p$-stable commutative semiring. Let $P$ be a $\text{\sf Datalog}^{\circ}$ program where the maximum number of multiplicands in any product is at most $\lambda$. Let $D$ be the input EDB database. Let $\sigma$ be number of the semiring elements referenced in $P$ or $D$. Let $n$ denote the total number of ground atoms in an IDB that at some point in the iterative evaluation of $P$ over semiring $\bm{S}$ on input $D$ have a nonzero associated semiring value. Then the iterative evaluation of $P$ over semiring $\bm{S}$ on input $D$ converges within
$\lceil pn(n+3)\cdot(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma+1)\rceil$
steps.
For a vector $\bm{v}\in\mathbb{N}^{\sigma}$, we let $Z(\bm{v})$ denote the product corresponding to $\bm{v}$, i.e. $\prod_{s=1}^{\sigma}a_{s}^{\bm{v}_{s}}$, where $a_{s}$ is the element corresponding to the $s$th entry of the vector. We naturally extend the notation to a vector set $\bm{V}$, by letting $Z(\bm{V}):=\bigoplus_{\bm{v}\in\bm{V}}Z(\bm{v})$. To prove the theorem, we need the following lemma, which roughly speaking shows that the summation of all products corresponding to vectors in $\mathcal{L}$ with coefficients up to $p$ doesn’t change when added a product corresponding to any other vector in $\mathcal{L}$. This lemma was proven in [7] (See Section 5.2, in particular the proof of Theorem 5.10 in the journal / ArXiV version of [7]) but we include the proof in the appendix for completeness.
Lemma 6.2.
Let $\mathcal{L}$ be a linear set with offset vector $\bm{v}_{0}$ and basis vectors $\bm{v}_{1},\bm{v}_{2},\ldots\bm{v}_{m}$. Let $\mathcal{L}_{\leq p}:=\{\bm{v}_{0}+\kappa_{1}\bm{v}_{1}+\kappa_{2}\bm{v}_{2}+\ldots+\kappa_{m}\bm{v}_{m}\;|\;\kappa_{i}\in[0,p]\forall\ i\}$. Consider an arbitrary $\bm{w}=\bm{v}_{0}+k_{1}\bm{v}_{1}+k_{2}\bm{v}_{2}+\ldots+k_{m}\bm{v}_{m}$ where $(k_{1},k_{2},\ldots,k_{m})\in\mathbb{N}^{m}$ and $k_{i}>p$ for some $i$. Then, we have
$Z(\mathcal{L}_{\leq p})=Z(\mathcal{L}_{\leq p})\oplus Z(\bm{w})$
We now have all tools to prove Theorem 1.1. Consider an arbitrary IDB variable $X_{r}$ and let $L$ be the CFL associated with this variable. Let $\mathcal{M}$ be a semi-linear set that satisfies the properties stated in Theorem 3.1. Let $\mathcal{T}^{r}_{q}$ denote the collection of the parse trees of depth at most $q$ starting with $X_{r}$. Our goal is to show:
(29) $f_{r}^{(q)}(\bm{0})=f_{r}^{(q+1)}(\bm{0})$
@@ 6. Bounding the Number of Iterations
$f_{r}^{(q)}(\bm{0})=\bigoplus_{T\in\mathcal{T}^{r}_{q}}Z(T),\textnormal{ and}$
(30) $q:=\lceil pn(n+3)\cdot(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma+1)\rceil$
Consider an arbitrary $T\in\mathcal{T}^{r}_{q+1}\setminus\mathcal{T}^{r}_{q}$. By Theorem 3.1 (3), $Y(T)$ has a linear representation
$\Psi(Y(T))=\bm{v}_{0}+k_{1}\bm{v}_{1}+k_{2}\bm{v}_{2}+\ldots+k_{m}\bm{v}_{m}$
within some $\mathcal{L}$ in $\mathcal{M}$ such that $1+k_{1}+k_{2}+\ldots+k_{m}>q/(n(n+3)/2)$. Thus, we have $k:=k_{1}+k_{2}+\ldots+k_{m}>ph$ where
$h:=2(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma)$
By Lemma 3.2, we can find a linear representation $\Psi(Y(T))=\bm{v}_{0}+k^{\prime}_{1}\bm{v}^{\prime}_{1}+k^{\prime}_{2}\bm{v}_{2}+\ldots+k^{\prime}_{h}\bm{v}^{\prime}_{h}$, where $k^{\prime}_{1}+k^{\prime}_{2}+\ldots+k^{\prime}_{h}>ph$. By the pigeonhole’s principle, we have that $k^{\prime}_{j}>p$ for some $j$.
Let $\mathcal{L}^{\prime}$ be the subset of $\mathcal{L}$ that only consists of basis vectors $\bm{v}^{\prime}_{1},\bm{v}^{\prime}_{2},\ldots,\bm{v}^{\prime}_{h}$ together with offset vector $\bm{v}_{0}$. Then, by Lemma 6.2, we have
$\bigoplus_{\bm{u}\in\mathcal{L}^{\prime}_{\leq p}}Z(\bm{u})=\bigoplus_{\bm{u}\in\mathcal{L}^{\prime}_{\leq p}}Z(\bm{u})\oplus Z(T),$
where we used $\bigoplus_{\bm{u}\in\mathcal{L}^{\prime}_{\leq p}}Z(\bm{u})=Z(\mathcal{L}^{\prime}_{\leq p})$, which is the case by definition. To complete the proof of Theorem 1.1, it is sufficient to show
(31) $\mathcal{L}^{\prime}_{\leq p}\subseteq\{\Psi(Y(T))\;|\;T\in\mathcal{T}_{q}^{r}\}$
To see this consider any $\bm{v}=\bm{v}_{0}+\kappa_{1}\bm{v}_{1}+\ldots+\kappa_{h}\bm{v}_{h}\in\mathcal{L}^{\prime}_{\leq p}$. By definition of $\mathcal{L}^{\prime}_{\leq p}$, $\kappa_{i}\leq p$ for all $i\in[h]$. Then, thanks to Theorem 3.1 property (2)(d), we know that there is a word $w\in L$ with $\Psi(w)=\bm{v}$ such that $w$ is generated by a parse tree $T^{\prime}$ of depth at most $(ph+1)n(n+3)/2\leq q$. Thus, it must be the case that $\bm{v}=Z(T^{\prime})\in\{Z(T)\;|\;T\in\mathcal{T}_{q}^{r}\}$. This establishes Eqn. (31) as desired, and therefore we have proven Theorem 1.1.
7. Conclusion
This paper considers the convergence of recursive $\text{\sf Datalog}^{\circ}$ programs using natural iterative evaluation over the semirings where convergence is not program dependent, namely the stable commutative semirings. Previously the best-known bound on convergence time was exponential in the output size. Our main contribution is to show that in fact the time to convergence can be bounded by a polynomial in the natural parameters, such as the output size. One consequence of this result is a better understanding of how much worse the time to convergence can be for general $\text{\sf Datalog}^{\circ}$ programs than linear $\text{\sf Datalog}^{\circ}$ programs. One reasonable interpretation of our results is that the worst-case time to convergence for general $\text{\sf Datalog}^{\circ}$ programs is not too much worse than the worst-case time to convergence for linear $\text{\sf Datalog}^{\circ}$ programs, which was a bit surprising to us given that generally one doesn’t expect algorithmic convergence bounds for non-linear optimization to be competitive with the bounds for linear optimization.
There are several natural directions for followup research. While essentially tight bounds are known for convergence time for linear $\text{\sf Datalog}^{\circ}$ programs, we do not establish the tightness of our bound. So one natural research direction is to determine tight bounds on the convergence rate for general $\text{\sf Datalog}^{\circ}$ programs. Another natural research direction is to show some sort of bounds on convergence time over non-stable semirings. Note that such bounds would have to be program-dependent. Another natural research direction would be to develop other algorithms for evaluating $\text{\sf Datalog}^{\circ}$ programs and analyze their convergence bounds.
Appendix A Omitted Proof
Proof of Lemma 6.2.
Assume wlog that $k_{1},\ldots,k_{m^{\prime}}>p$ and $k_{m^{\prime}+1},\ldots,k_{m}\leq p$. Let $G_{j}:=\{\bm{\kappa}=(\kappa_{1},\ldots,\kappa_{m}\;|\;\kappa_{i}\in[0,k_{i}]\ \forall i\in[0,j]\cup[m^{\prime}+1,m]\textnormal{ and }\kappa_{i}\in[0,p]\ \forall i\in[j+1,m^{\prime}]\}$ for all $j\in[0,m^{\prime}]$. Note that to prove the lemma it suffices to show
$Z(G_{0})=Z(G_{0})\oplus Z(\bm{w})$
because $\{\bm{v}_{0}+\kappa_{1}\bm{v}_{1}+\ldots+\kappa_{m}\bm{v}_{m}\;|\;\bm{\kappa}\in G_{0}\}\subseteq\mathcal{L}_{\leq p}$. We are going to establish
(32) $Z(G_{0})=Z(G_{1})=\ldots=Z(G_{m^{\prime}})$
and
(33) $Z(G_{j})\oplus Z(G_{j+1}\setminus G_{j})=Z(G_{j})\;\forall j\in[0,m^{\prime}-1]$
Indeed if we have them,
$Z(G_{0})\oplus Z(\bm{w})=Z(G_{m^{\prime}-1})\oplus Z(\bm{w})=Z(G_{m^{\prime}-1})=Z(G_{0}),$
as desired, since $(k_{1},k_{2},\ldots,k_{m})\in G_{m^{\prime}}\setminus G_{m^{\prime}-1}$ and $\bm{w}=\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{m}\bm{v}_{m}$.
It now remains to show Eqn.  (32) and (33). Consider a fixed $j\in[0,m^{\prime}-1]$. Consider an arbitrary $\bm{\kappa}\in G_{j}$. Let $\bm{\kappa}^{\prime}(q)$ be $\bm{\kappa}$ with $\bm{\kappa}_{j+1}$ ($j+1$-th coordinate of $\bm{\kappa}$) replaced with $q$. Let $\bm{v}(\bm{\kappa}^{\prime}(q)):=\bm{v}_{0}+\kappa^{\prime}_{1}(q)\bm{v}_{1}+\ldots+\kappa^{\prime}_{m}(q)\bm{v}_{m}$ be the vector represented by the linear representation $\bm{\kappa}^{\prime}(q)$. Let $z_{i}:=Z(\bm{v}_{i})$. Then, we have
$\bigoplus_{q=0}^{p}Z(\bm{v}(\bm{\kappa}^{\prime}(q)))$ $=\bigoplus_{q=0}^{p}\left(\prod_{i=1:i\neq j+1}^{m}z_{i}^{\kappa_{i}}\right)z_{j+1}^{q}=\prod_{i=1:i\neq j+1}^{m}z_{i}^{\kappa_{i}}\bigoplus_{q=0}^{p}z_{j+1}^{q}$
$=\prod_{i=1:i\neq j+1}^{m}z_{i}^{\kappa_{i}}z_{j+1}^{(p)}$
$=\prod_{i=1:i\neq j+1}^{m}z_{i}^{\kappa_{i}}z_{j+1}^{(k_{j+1})}\quad\mbox{[$p$-stability and $k_{j+1}>p$]}$
$=\bigoplus_{q=0}^{p}\left(\prod_{i=1:i\neq j+1}^{m}z_{i}^{\kappa_{i}}\right)z_{j+1}^{q}+\bigoplus_{q=p+1}^{k_{j+1}}\left(\prod_{i=1:i\neq j+1}^{m}z_{i}^{\kappa_{i}}\right)z_{j+1}^{q}$
$=\bigoplus_{q=0}^{p}Z(\bm{v}(\bm{\kappa}^{\prime}(q)))+\bigoplus_{q=p+1}^{k_{j+1}}Z(\bm{v}(\bm{\kappa}^{\prime}(q)))$
Since $\bigcup_{q=0}^{p}\bm{v}(\bm{\kappa}^{\prime}(q))\subseteq G_{j}$, and any vector in $G_{j+1}\setminus G_{j}$ is of the form of $\bm{\kappa}^{\prime}(q)$ for some $q\in[p+1,k_{j+1}]$ for some $\bm{\kappa}\in G_{j}$, we have $Z(G_{j})=Z(G_{j+1})$. In other words, we showed that any product corresponding to $G_{j+1}\setminus G_{j}$ is subsumed by some $p+1$ products in $Z(G_{j})$ using the $p$-stability. For the same reason, we have Eqn. (33). ∎